OpenStudy (anonymous):

guys help me please im in danger!! . Find the slope of the line that passes through the points (3, -6) and (6, 12).

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

ok dnt leave me stay here wen im done solving it

OpenStudy (anonymous):

slope of a line: (y1-y2)/(x1-x2), plug in the values

OpenStudy (anonymous):

m=(12-(-6))/(6-3) = 18/3 = 6

OpenStudy (anonymous):

@privetek is it y2-y1 or y1-y2? look at the way @emilie.li told me

OpenStudy (anonymous):

it's (y2-y1)/(x2-x1 )

OpenStudy (anonymous):

it doesn't matter if it's y1-y2 or y2-y1 as long as you keep it consistent for both the x and y values.

OpenStudy (anonymous):

yes, i guess @emilie.li is right.. :)

OpenStudy (anonymous):

that means that you can't write (y1-y2)/(x2-x1) because you won't get the right answer for the slope

OpenStudy (anonymous):

so @emilie.li @privetek answer was wrong?

OpenStudy (anonymous):

your answer is right, I was just saying that you have to be consistent with how you subtract the values for both x and y.